OpenStudy (anonymous):

A rectangular area of 3750 square feet is to be fenced off. Two opposite sides will have fencing that costs $2 per foot and the remaining sides will ise fencing that costs $3 per foot. Find the dimensions of the rectangle that will minimize the cost.

OpenStudy (anonymous):



OpenStudy (anonymous):

\(C=2x+2x+3y+3y=4x+6y\) and \(xy=3750\) so \(y=\frac{3750}{x}\) and now \[C(x)=4x+6\times \frac{3750}{x}\]

OpenStudy (anonymous):

minimize that by taking the derivative and finding the critical points for \(x>0\)

OpenStudy (anonymous):

critical points would be 4

OpenStudy (anonymous):

total answer for 5641

OpenStudy (anonymous):

seems unlikely

OpenStudy (anonymous):

what did I do wrong

OpenStudy (anonymous):

\[C(x)=4x+\frac{22500}{x}\] \[C'(x)=4-\frac{22500}{x^2}\] set this equal to zero and solve

OpenStudy (anonymous):

\[4-\frac{22500}{x^2}=0\] \[\frac{22500}{x^2}=4\] \[4x^2=22500\] etc

OpenStudy (anonymous):

i think you get \(x=75\)

OpenStudy (anonymous):

ok I get 75 and -75

OpenStudy (anonymous):

forget the negative one

OpenStudy (anonymous):

ok so then 75 is that it

OpenStudy (anonymous):

is that all that I need